# MEGADance: Mixture-of-Experts Architecture for Genre-Aware 3D Dance Generation

### v2

@@ 3.3.3 Mamba-Transformer Hybrid Backbone
Cross-Modal Global-Context Modeling. Leveraging the Transformer’s global receptive field vaswani2017attention , we concatenate cross-modal features along the temporal axis and facilitate structured interactions among music, upper-body, and lower-body representations, by a Attention layer and a Feed Forward layer. The attention layer is the core component that defines the computational dependencies among sequential data elements and is implemented as:
@@
where $Q,K,V$ denote the query, key, and value from input, and $M$ is the mask, which determines the type of attention layers. The two most common attention types are full attention vaswani2017attention , which enables global context exchange, and causal attention radford2018improving , which restricts each position to attend only to current and past inputs. Given that music-to-dance generation is typically applied to long sequence while being constrained by limited computational resources, training is commonly conducted on short clips. During inference, the sequence is first extended autoregressively up to the training length (step 1), and then completed using a sliding window approach for the remaining part (step 2). Training driven by standard causal attention only aligns with step 1 and fails to account for the dominant step 2 during inference, thereby limiting generation performance. To better align training with inference, we introduce a sliding-window attention mechanism that mimics the generation process. The attention mask $M\in\mathbb{R}^{3T\times 3T}$ is structured as a $3\times 3$ block matrix, where each block is a $T\times T$ sliding-window mask, enabling cross-modal global-context attention.
Intra-Modal Local-Dependency Modeling. While the Transformer excels at temporal modeling, it is inherently position-invariant and captures sequence order only through positional encodings vaswani2017attention , which limits its deep understanding of local dependencies. In contrast, music-to-dance generation demands strong local continuity between movements. Owing to its inherent sequential inductive bias, Mamba gu2023mamba has demonstrated strong performance in modeling fine-grained local dependencies xu2024mambatalk ; fu2024mambagesture . We therefore apply independent Mamba to the music, upper-body, and lower-body features respectively, to model their intra-modal local dependencies. Specifically, Selective State Space model (Mamba) incorporates a Selection mechanism and a Scan module (S6) gu2023mamba to dynamically select salient input segments for efficient sequence modeling. Unlike the traditional S4 model gu2021efficiently with fixed parameters $A,B,C$, and scalar $\Delta$, Mamba adaptively learns these parameters via fully-connected layers, enhancing generalization capabilities. Mamba employs structured state-space matrices, imposing constraints that improve computational efficiency. For each time step $t$, the input $x_{t}$, hidden state $h_{t}$, and output $y_{t}$ follow:
@@ 4.1 Dataset
FineDance. FineDance li2023finedance is the largest public dataset for 3D music-to-dance generation, featuring professionally performed dances captured via optical motion capture. It provides 7.7 hours of motion data at 30 fps across 16 distinct dance genres. Following li2024lodge , we evaluate on 20 test-set music clips, generating 1024-frame (34.13s) dance sequences.
AIST++. AIST++ li2021ai is a widely used benchmark comprising 5.2 hours of 60 fps street dance motion, covering 10 dance genres. Following li2021ai , we use 40 test-set music clips to generate 1200-frame (20.00s) sequences.
@@ 4.2 Quantitative Evaluation
Bailando++siyao2023bailando++ 54.79 16.29 8.42 6.18 5.98 4.73 0.213 3.85 3.50 3.82
FineNetli2023finedance 65.15 23.81 13.22 5.84 5.19 4.29 0.219 3.62 3.65 3.47
Lodgeli2024lodge 55.03 14.87 5.22 6.14 6.18 5.50 0.218 4.18 4.17 4.08
@@
Comparison. We evaluate MEGADance against state-of-the-art (SOTA) baselines on both the FineDance and AIST++ datasets using a comprehensive suite of metrics. For each generated sequence, we compute Fréchet Inception Distance ($FID$) and Diversity ($DIV$) across three feature spaces: (1) Kinetic (k), capturing motion dynamics; (2) Geometric (g), encoding spatial joint relations; and (3) Style (s), extracted via a Transformer-based genre classifier. We also assess music-motion synchronization using Beat Align Score ($BAS$), following li2021ai ; li2024lodge . On the FineDance dataset (Tab. 1), MEGADance outperforms all baselines, achieving the lowest FID in all three feature types ($FID_{k}$=50.00, $FID_{g}$=13.02, $FID_{s}$=2.52), the highest motion diversity ($DIV_{k}$=6.23, $DIV_{g}$=6.27, $DIV_{s}$=5.78), and the best BAS (0.226). On the AIST++ datasets (Tab. 4.3), MEGADance again ranks first in FID ($FID_{k}$=25.89, $FID_{g}$=12.62), and achieves strong performance in diversity ($DIV_{g}$=5.84, $DIV_{s}$=6.23) and BAS (0.238). These results underscore the effectiveness of our genre-aware Mixture-of-Experts design in balancing motion quality, creativity, and synchronization across diverse datasets.
User Study. Dance’s inherent subjectivity makes user feedback essential for evaluating generated movementslegrand2009perceiving in the music-to-dance generation task. We select 30 in-the-wild music segments (34 seconds each) and generate dance sequence using above models. These sequences are evaluated through a double-blind questionnaire, by 30 participants with backgrounds in dance practice, including undergraduate and graduate-level students. The questionnaires are based on a 5-point scale (Great, Good, Fair, Bad, Terrible) and assess three aspects: Dance Synchronization ($DS$, alignment with rhythm and style), Dance Quality ($DQ$, biomechanical plausibility and aesthetics), and Dance Creativity ($DC$, originality and range). As shown in Tab. 1, MEGADance significantly outperforms all baselines across user-rated metrics (i.e., $DS=4.30$, $DQ=4.25$, $DC=4.23$). Its high scores in various aspects underscore its superiority in generating movements in terms of human preferences.
@@ 4.3 Qualitative Evaluation
To assess the visual quality of the generated dance sequences, we perform a qualitative comparison between MEGADance and several existing baseline models, as depicted in Fig. 3. In terms of expressiveness, MEGADance outperforms the competing methods in several key areas. For instance, Lodgeli2024lodge struggles with stylistic consistency, often blending conflicting dance genres, such as incorporating Uyghur movements into a typical Breaking routine, leading to a disjointed aesthetic. FineNetli2023finedance , while capable of generating movement sequences, suffers from significant artifacts, including unnatural sliding and teleportation, which detract from the fluidity and physical realism of the motions. Additionally, Bailando++siyao2023bailando++ demonstrates a lack of diversity, with movements frequently repeating and a heavy reliance on static, in-place gestures, limiting the range of expressive movement patterns. These findings underscore the superiority of MEGADance in generating diverse, genre-consistent, and musically synchronized dance sequences.
@@
FACTli2021ai 35.35 22.11 5.94 6.18 0.221
Bailando++siyao2023bailando++ 30.21 15.48 5.35 5.13 0.228
EDGEtseng2023edge 42.16 22.12 3.96 4.61 0.233
Lodgeli2024lodge 35.72 17.92 5.72 5.91 0.247
@@
FineNetli2023finedance 13.22 4.29 42.06 37.44
Lodgeli2024lodge 5.22 5.50 51.86 45.23
@@ 4.4 Genre Controllability Evaluation
To quantitatively assess genre controllability, we compare MEGADance with Lodge li2024lodge and FineNet li2023finedance . Using ground-truth genre labels, we evaluate style alignment ($FID_{s}$) and style diversity ($DIV_{s}$) on 20 test clips. We further assess genre classification accuracy ($ACC$) and F1 score ($F1$), conditioned on the correct genre and four randomly sampled negative genres. As shown in Tab. 4.3, MEGADance achieves the best performance across all metrics. It significantly reduces $FID_{s}$ (2.52) while improving diversity ($DIV_{s}$ = 5.78). Despite potential cross-modal conflicts (e.g., assigning a Popping genre to a typical Chinese music clip), MEGADance achieves high genre discriminability ($ACC$ = 75.64, $F1$ = 70.81), closely approaching ground-truth performance. Compared to FineNet and Lodge, our method produces motion sequences that are both more stylistically coherent and genre-distinctive. Our MoE-based genre routing prevents cross-genre interference via disentangled expert subspaces activated by discrete labels, whereas naive continuous fusion (e.g. Feature Addition in li2023finedance ; li2024lodge or Cross Attention in huang2022genre ) inherently blurs stylistic boundaries.
@@ 4.5.2 High-Fidelity Dance Quantization Stage
We investigate the effectiveness of three key components in High-Fidelity Dance Quantization Stage: Finite Scalar Quantization (FSQ), the Kinematic Loss ($\mathcal{L}_{\text{Kin.}}$), and the Dynamic Loss ($\mathcal{L}_{\text{Dyn.}}$). Tab. 4(b) reports the $MSE$ and $MAE$ on both SMPL parameters and 3D joint positions. (1) FSQ. Replacing FSQ with VQ-VAE leads to a significant performance drop in all metrics (e.g., Joint $MSE$ increases from 0.0069 to 0.0220), validating FSQ’s superiority. Moreover, replacing VQ-VAE with FSQ achieves full codebook utilization (75% → 100%). (2) Kinematic Loss. Removing the kinematic loss notably increases errors in both SMPL ($MSE$: $0$ 0.0238) and joint space ($MAE$: 0.0469 → 0.0507), highlighting its role in enforcing accurate structural constraints via forward kinematics. (3) Dynamic Loss. Excluding the dynamic loss results in a moderate degradation in temporal fidelity (e.g., Joint $MSE$: 0.0069 → 0.0073), demonstrating its contribution for temporal fidelity.

### v3

@@ 3.3.3 Mamba-Transformer Hybrid Backbone
Cross-Modal Global-Context Modeling. Leveraging the Transformer’s global receptive field Vaswani (2017), we concatenate cross-modal features along the temporal axis and facilitate structured interactions among music, upper-body, and lower-body representations, by a Attention layer and a Feed Forward layer. The attention layer is the core component that defines the computational dependencies among sequential data elements and is implemented as:
@@
where $Q,K,V$ denote the query, key, and value from input, and $M$ is the mask, which determines the type of attention layers. The two most common attention types are full attention Vaswani (2017), which enables global context exchange, and causal attention Radford et al. (2018), which restricts each position to attend only to current and past inputs. Given that music-to-dance generation is typically applied to long sequence while being constrained by limited computational resources, training is commonly conducted on short clips. During inference, the sequence is first extended autoregressively up to the training length (step 1), and then completed using a sliding window approach for the remaining part (step 2). Training driven by standard causal attention only aligns with step 1 and fails to account for the dominant step 2 during inference, thereby limiting generation performance. To better align training with inference, we introduce a sliding-window attention mechanism that mimics the generation process. The attention mask $M\in\mathbb{R}^{3T\times 3T}$ is structured as a $3\times 3$ block matrix, where each block is a $T\times T$ sliding-window mask, enabling cross-modal global-context attention.
Intra-Modal Local-Dependency Modeling. While the Transformer excels at temporal modeling, it is inherently position-invariant and captures sequence order only through positional encodings Vaswani (2017), which limits its deep understanding of local dependencies. In contrast, music-to-dance generation demands strong local continuity between movements. Owing to its inherent sequential inductive bias, Mamba Gu and Dao (2023) has demonstrated strong performance in modeling fine-grained local dependencies Xu et al. (2024); Fu et al. (2024). We therefore apply independent Mamba to the music, upper-body, and lower-body features respectively, to model their intra-modal local dependencies. Specifically, Selective State Space model (Mamba) incorporates a Selection mechanism and a Scan module (S6) Gu and Dao (2023) to dynamically select salient input segments for efficient sequence modeling. Unlike the traditional S4 model Gu et al. (2021) with fixed parameters $A,B,C$, and scalar $\Delta$, Mamba adaptively learns these parameters via fully-connected layers, enhancing generalization capabilities. Mamba employs structured state-space matrices, imposing constraints that improve computational efficiency. For each time step $t$, the input $x_{t}$, hidden state $h_{t}$, and output $y_{t}$ follow:
@@ 4.1 Dataset
FineDance. FineDance Li et al. (2023) is the largest public dataset for 3D music-to-dance generation, featuring professionally performed dances captured via optical motion capture. It provides 7.7 hours of motion data at 30 fps across 16 distinct dance genres. Following Li et al. (2024c), we evaluate on 20 test-set music clips, generating 1024-frame (34.13s) dance sequences.
AIST++. AIST++ Li et al. (2021) is a widely used benchmark comprising 5.2 hours of 60 fps street dance motion, covering 10 dance genres. Following Li et al. (2021), we use 40 test-set music clips to generate 1200-frame (20.00s) sequences.
@@ 4.2 Quantitative Evaluation
Bailando++Siyao et al. (2023) 54.79 16.29 8.42 6.18 5.98 4.73 0.213 3.85 3.50 3.82
FineNetLi et al. (2023) 65.15 23.81 13.22 5.84 5.19 4.29 0.219 3.62 3.65 3.47
LodgeLi et al. (2024c) 55.03 14.87 5.22 6.14 6.18 5.50 0.218 4.18 4.17 4.08
@@
Comparison. We evaluate MEGADance against state-of-the-art (SOTA) baselines on both the FineDance and AIST++ datasets using a comprehensive suite of metrics. For each generated sequence, we compute Fréchet Inception Distance ($FID$) and Diversity ($DIV$) across three feature spaces: (1) Kinetic (k), capturing motion dynamics; (2) Geometric (g), encoding spatial joint relations; and (3) Style (s), extracted via a Transformer-based genre classifier. We also assess music-motion synchronization using Beat Align Score ($BAS$), following Li et al. (2021, 2024c). On the FineDance dataset (Tab. 1), MEGADance outperforms all baselines, achieving the lowest FID in all three feature types ($FID_{k}$=50.00, $FID_{g}$=13.02, $FID_{s}$=2.52), the highest motion diversity ($DIV_{k}$=6.23, $DIV_{g}$=6.27, $DIV_{s}$=5.78), and the best BAS (0.226). On the AIST++ datasets (Tab. 4.3), MEGADance again ranks first in FID ($FID_{k}$=25.89, $FID_{g}$=12.62), and achieves strong performance in diversity ($DIV_{g}$=5.84, $DIV_{s}$=6.23) and BAS (0.238). These results underscore the effectiveness of our genre-aware Mixture-of-Experts design in balancing motion quality, creativity, and synchronization across diverse datasets.
User Study. Dance’s inherent subjectivity makes user feedback essential for evaluating generated movementsLegrand and Ravn (2009) in the music-to-dance generation task. We select 30 in-the-wild music segments (34 seconds each) and generate dance sequence using above models. These sequences are evaluated through a double-blind questionnaire, by 30 participants with backgrounds in dance practice, including undergraduate and graduate-level students. The questionnaires are based on a 5-point scale (Great, Good, Fair, Bad, Terrible) and assess three aspects: Dance Synchronization ($DS$, alignment with rhythm and style), Dance Quality ($DQ$, biomechanical plausibility and aesthetics), and Dance Creativity ($DC$, originality and range). As shown in Tab. 1, MEGADance significantly outperforms all baselines across user-rated metrics (i.e., $DS=4.30$, $DQ=4.25$, $DC=4.23$). Its high scores in various aspects underscore its superiority in generating movements in terms of human preferences.
@@ 4.3 Qualitative Evaluation
To assess the visual quality of the generated dance sequences, we perform a qualitative comparison between MEGADance and several existing baseline models, as depicted in Fig. 3. In terms of expressiveness, MEGADance outperforms the competing methods in several key areas. For instance, LodgeLi et al. (2024c) struggles with stylistic consistency, often blending conflicting dance genres, such as incorporating Uyghur movements into a typical Breaking routine, leading to a disjointed aesthetic. FineNetLi et al. (2023), while capable of generating movement sequences, suffers from significant artifacts, including unnatural sliding and teleportation, which detract from the fluidity and physical realism of the motions. Additionally, Bailando++Siyao et al. (2023) demonstrates a lack of diversity, with movements frequently repeating and a heavy reliance on static, in-place gestures, limiting the range of expressive movement patterns. These findings underscore the superiority of MEGADance in generating diverse, genre-consistent, and musically synchronized dance sequences.
@@
FACTLi et al. (2021) 35.35 22.11 5.94 6.18 0.221
Bailando++Siyao et al. (2023) 30.21 15.48 5.35 5.13 0.228
EDGETseng et al. (2023) 42.16 22.12 3.96 4.61 0.233
LodgeLi et al. (2024c) 35.72 17.92 5.72 5.91 0.247
@@
FineNetLi et al. (2023) 13.22 4.29 42.06 37.44
LodgeLi et al. (2024c) 5.22 5.50 51.86 45.23
@@ 4.4 Genre Controllability Evaluation
To quantitatively assess genre controllability, we compare MEGADance with Lodge Li et al. (2024c) and FineNet Li et al. (2023). Using ground-truth genre labels, we evaluate style alignment ($FID_{s}$) and style diversity ($DIV_{s}$) on 20 test clips. We further assess genre classification accuracy ($ACC$) and F1 score ($F1$), conditioned on the correct genre and four randomly sampled negative genres. As shown in Tab. 4.3, MEGADance achieves the best performance across all metrics. It significantly reduces $FID_{s}$ (2.52) while improving diversity ($DIV_{s}$ = 5.78). Despite potential cross-modal conflicts (e.g., assigning a Popping genre to a typical Chinese music clip), MEGADance achieves high genre discriminability ($ACC$ = 75.64, $F1$ = 70.81), closely approaching ground-truth performance. Compared to FineNet and Lodge, our method produces motion sequences that are both more stylistically coherent and genre-distinctive. Our MoE-based genre routing prevents cross-genre interference via disentangled expert subspaces activated by discrete labels, whereas naive continuous fusion (e.g. Feature Addition in Li et al. (2023, 2024c) or Cross Attention in Huang et al. (2022)) inherently blurs stylistic boundaries.
@@ 4.5.2 High-Fidelity Dance Quantization Stage
We investigate the effectiveness of three key components in High-Fidelity Dance Quantization Stage: Finite Scalar Quantization (FSQ), the Kinematic Loss ($\mathcal{L}_{\text{Kin.}}$), and the Dynamic Loss ($\mathcal{L}_{\text{Dyn.}}$). Tab. 4(b) reports the $MSE$ and $MAE$ on both SMPL parameters and 3D joint positions. (1) FSQ. Replacing FSQ with VQ-VAE leads to a significant performance drop in all metrics (e.g., Joint $MSE$ increases from 0.0069 to 0.0220), validating FSQ’s superiority. Moreover, replacing VQ-VAE with FSQ achieves full codebook utilization (75% → 100%). (2) Kinematic Loss. Removing the kinematic loss notably increases errors in both SMPL ($MSE$: 0.0200 → 0.0238) and joint space ($MAE$: 0.0469 → 0.0507), highlighting its role in enforcing accurate structural constraints via forward kinematics. (3) Dynamic Loss. Excluding the dynamic loss results in a moderate degradation in temporal fidelity (e.g., Joint $MSE$: 0.0069 → 0.0073), demonstrating its contribution for temporal fidelity.